Calculate the partition function, free energy, entropy, Cv and Cp of N linear

harmonic oscillators.


1
Expert's answer
2021-03-16T08:44:22-0400
"\varDelta Z=\int_1^n\frac {dU+\delta A}{T}"

"dU=\frac {m}{\mu}{i}{2}RdT"

"dS=nRdlnV"

"c_V=\frac {i}{2}R"


"c_P=\frac {i+2}{2}R"